### v1

@@ Abstract
Accurately and swiftly assessing damage from conflicts is crucial for humanitarian aid and regional stability. In conflict zones, damaged zones often share similar architectural styles, with damage typically covering small areas and exhibiting blurred boundaries. These characteristics lead to limited data, annotation difficulties, and significant recognition challenges, including high intra-class similarity and ambiguous semantic changes. To address these issues, we introduce a pre-trained DINOv3 model and propose a multi-scale cross-attention difference siamese network (MC-DiSNet). The powerful visual representation capability of the DINOv3 backbone enables robust and rich feature extraction from bi-temporal remote sensing images. The multi-scale cross-attention mechanism allows for precise localization of subtle semantic changes, while the difference siamese structure enhances inter-class feature discrimination, enabling fine-grained semantic change detection. Furthermore, a simple yet powerful lightweight decoder is designed to generate clear detection maps while maintaining high efficiency. We also release a new Gaza-change dataset containing high-resolution satellite image pairs from 2023-2024 with pixel-level semantic change annotations. It is worth emphasizing that our annotations only include semantic pixels of changed areas. Unlike conventional semantic change detection (SCD), our approach eliminates the need for large-scale semantic annotations of bi-temporal images, instead focusing directly on the changed regions. We term this new task change semantic detection (CSD). The CSD task represents a direct extension of binary change detection (BCD). Due to the limited spatial extent of semantic regions, it presents greater challenges than traditional SCD tasks. We evaluated our method under the CSD framework on both the Gaza-Change and SECOND datasets. Experimental results demonstrate that our proposed approach effectively addresses the CSD task, and its outstanding performance paves the way for practical applications in rapid damage assessment across conflict zones.
@@ 1 Introduction
To address the aforementioned challenges, we propose a new paradigm termed change semantic detection (CSD). Distinct from conventional semantic change detection (SCD), our approach eliminates the need for annotating entire semantic regions, instead focusing solely on change masks. This framework represents a direct extension of binary change detection (BCD). While significantly reducing annotation difficulty and time requirements, this new paradigm consequently increases the challenge of limited target region proportions. This necessitates substantial improvements in the model’s capability to extract features from small target areas.
@@
Figure 1: Evolution of change detection paradigms: (a) Binary Change Detection (BCD), (b) Semantic Change Detection (SCD), and (c) our proposed Change Semantic Detection (CSD).
@@
Based on the above analysis, we identify four major challenges in framing conflict-induced damage assessment as a CSD task: (1) Inherent data scarcity: Limited by the geographical extent of conflict zones and the number of destroyed areas available for training. (2) Small target regions: CSD focuses exclusively on damaged areas, resulting in minimal semantic region coverage. (3) Subtle and ambiguous changes: Infrastructure damage in conflict zones varies significantly in severity and extent, particularly making minor damage difficult to detect. (4) High inter-class similarity: Different facility categories within the same region may share similar characteristics, making fine-grained damage assessment particularly challenging for semantic change detection.
@@
To bridge this gap, we draw inspiration from the recent success of foundational vision models. We argue that leveraging their rich, pre-trained representations is key to overcoming data scarcity and recognizing subtle semantic changes. In this paper, we propose a novel DINOv3-driven siamese network for CSD. Specifically, we adopt the DINOv3 simeoni2025dinov3 model pre-trained on satellite data, with ConvNeXt liu2022convnet as its main backbone architecture, which helps reduce the distribution discrepancy between the pre-training data and the actual application data. Then, we propose a multi-scale attention mechanism to extract and enhance features at different levels, aiming to capture the subtle and ambiguous change features of infrastructure damage. Furthermore, we perform an absolute value differential operation on the obtained semantic-rich feature maps to increase inter-class feature differences. Finally, a carefully designed decoder network with attention enhancement is used to generate clear semantic change detection maps. We also release a building semantic change detection dataset of the Gaza area from 2023 to 2024. As shown in Figure 2, we present panoramic remote sensing images of the Gaza Strip captured by satellites. To the best of our knowledge, this is the first remote sensing semantic change detection study focused on conflict area assessment, laying a foundation for future research in related fields. In summary, our work makes the following key contributions:
@@
  • We propose the novel change semantic detection (CSD) paradigm that fundamentally shifts from exhaustive bi-temporal semantic annotation to focused labeling of changed semantic regions. This strategic simplification significantly reduces annotation complexity and human labor.
  • Extensive experiments show that our method achieves state-of-the-art performance not only on our proposed Gaza dataset but also on the public benchmarks, SECOND yang2021asymmetric, demonstrating its superior robustness and generalization capability.
@@ 3.2 Enhanced DINOv3 Backbone
To maximize the feature extraction capability of the DINOv3 backbone while balancing computational cost and accuracy for the proposed change semantic detection (CSD) task, we introduce an enhanced DINOv3 encoder. This encoder keeps the original ConvNeXt-Tiny weights frozen and enriches multi-scale features C3-C5 through three lightweight plug-ins. First, a bottleneck adapter is inserted after each MLP block, comprising a down-projection, GELU activation lee2023mathematical, and an up-projection branch with a reduction ratio. Its forward pass is formulated as:
@@ 3.3 Multi-Scale Cross-Attention Difference Module
Next, the concatenated tensor $\mathbf{F}_{\mathrm{cat}}$ == $\mathbf{C}_{\mathrm{fused}}^{t_{1}};\mathbf{C}_{\mathrm{fused}}^{t_{2}}]$ is fed to a lightweight stack of $3\times 3$ conv, BN, GELU and $1\times 1$ conv to model non-linear temporal discrepancies and produce an adaptive difference feature $\mathbf{D}_{\mathrm{ada}}$.
Finally, the direct difference $\mathbf{D}_{\mathrm{dir}}=|\mathbf{C}^{t_{1}}-\mathbf{C}^{t_{2}}|$ and the adaptive difference $\mathbf{D}_{\mathrm{ada}}$ are concatenated and compressed by $1\times 1$ conv, BN and GELU, yielding the multi-scale change feature:
@@
3.4 Context-Aware Feature Enhancement Decoder
To refine the coarse change cues produced by the MSCAD module, we design a context-aware feature enhancement decoder, which contains two lightweight yet effective sub-blocks. First, a three-layer depthwise-separable residual stack (Figure 3) enriches spatial context via
@@ 3.4 Context-Aware Feature Enhancement Decoder
$\mathbf{D}_{\mathrm{out}}=\mathbf{G}\odot\mathbf{Y},$ (8)
@@ 4.1.1 Datasets Description
Table 1 summarizes the specific information of two aforementioned datasets. It is important to emphasize that while both the Gaza-Change and SECOND datasets are annotated with semantic labels. To validate our proposed task of semantic change detection, the labels in the SECOND dataset were also processed to retain semantic information exclusively for the changed areas.
Table 1: Main information of two datasets.
@@ 4.1.3 Experimental Details
4.2 Results
We report the results achieved by the model checkpoint that performs best on the test set. Table 2 summarizes the quantitative comparison between MC-DiSNet and nine representative methods, including both CNN-based (SNUNet fang2021snunet) and Transformer-based (BIT chen2021remote, ChangeFormer bandara2022transformer) baselines, as well as the most recent ChangeMamba chen2024changemamba.
@@ 4.2 Results
Method Params (M) P(%) R(%) mIoU(%) F1(%) Best Epoch
USSFCNET 1.50 46.66 48.75 34.43 47.20 51
BIT 11.90 57.45 54.90 42.13 56.15 65
SNUNet 12.00 51.44 51.42 36.50 50.35 49
DDLNET 13.80 56.61 59.96 42.23 57.22 34
STNet 14.60 59.15 54.96 42.81 56.64 53
CDxLstm 16.20 56.96 47.29 36.84 51.25 83
ChangeFormer 41.10 59.06 46.52 36.62 50.50 88
Rsmamba 52.00 58.43 42.81 36.45 49.40 66
LGPNet 71.00 55.69 55.69 48.64 54.64 64
Ours 0.97 86.10 60.98 55.16 69.25 29
To further verify the effectiveness and generalizability of the proposed method, we conduct experiments on an extra SCD datasets: SECOND. Quantitative results are summarised in Table 3. Obviously, our method achieves an mIoU of 27.74% and an F1-score of 40.59%, surpassing the strongest competitor (BIT) by 6.87 percentage points in mIoU and 8.39 percentage points in F1-score. This consistent superiority on an external dataset demonstrates that the proposed modules do not overfit the primary training data and generalize effectively to other remote sensing scenarios with varying imaging conditions and annotation granularity.
Table 3: Performance comparison on SECOND datasets.
Method SECOND
mIoU (%) F1 (%)
BIT 20.87 32.20
SNUNet 12.88 20.85
SARASNet 16.38 25.34
USSFCNet 12.03 17.80
ChangeFormer 15.48 28.30
CDxLstm 15.19 27.25
RSMamba 12.54 20.09
LGPNet 12.42 19.37
Ours 27.74 40.59
As demonstrated on our own dataset as well as on the additional benchmarks, the proposed approach consistently delivers state-of-the-art results not only on small-scale sets (hundreds of image pairs) but also on large-scale collections (tens of thousands of samples). This robust performance across dramatically different data volumes indicates strong generalization capability and practical deployment potential for diverse remote-sensing scenarios.
4.3 Ablation Study
4.3.1 Module-level Ablation
MSCAD fuses multi-scale C3-C5 features into a shared 256-D space and re-weights them with a channel-wise gate, allowing the model to exploit both fine detail and high-level semantics. We evaluated the multi-scale attention mechanism in MSCAD, as shown in Table 4.
@@ 4.3.1 Module-level Ablation
Module P(%) R(%) mIoU (%) F1 (%)
w/ attention 86.10 60.98 55.16 69.25
w/o attention 89.37 54.35 50.28 64.99
Compared with the baseline without attention, the method incorporating attention mechanism achieved a 4.88% improvement in mIoU and a 4.26% increase in F1 score. These results demonstrate that the multi-scale attention mechanism significantly enhances the model’s feature extraction capability. Although precision shows a slight decrease, the substantial improvements in both recall and F1 score confirm the effectiveness of multi-scale features in detecting small targets.
4.3.2 Parameter-level Ablation
Table 5 compares different combinations of rank $r$ and scaling factor $\alpha$. As shown in Table.5, the best trade-off between accuracy and parameter budget is achieved at r=24 and $\alpha$=48, where the model reaches 69.25% F1 and 55.16% mIoU with only 0.968 M additional trainable weights.
@@ 4.3.2 Parameter-level Ablation
5 Discussion
Second, we propose a new paradigm that directly annotates changed areas rather than separately labeling bi-temporal semantic regions. This approach, which we term Change Semantic Detection (CSD), effectively extends traditional Binary Change Detection (BCD) to semantic change analysis while avoiding error accumulation from intermediate processing stages. The CSD framework substantially reduces both annotation workload and complexity, enabling rapid damage assessment in conflict scenarios.
Table 6: Quantitative results of our MC-DiSNet on the proposed Gaza-change dataset under the CSD task, reporting per-category performance metrics for all six change types.
@@ 5 Discussion
These findings highlight both the promise and challenges of automated damage assessment in conflict zones, while demonstrating the potential of our CSD paradigm and MC-DiSNet architecture to advance the field of humanitarian remote sensing.

### v2

@@ Abstract
Accurately and swiftly assessing damage from conflicts is crucial for humanitarian aid and regional stability. In conflict zones, damaged zones often share similar architectural styles, with damage typically covering small areas and exhibiting blurred boundaries. These characteristics lead to limited data, annotation difficulties, and significant recognition challenges, including high intra-class similarity and ambiguous semantic changes. To address these issues, we introduce a pre-trained DINOv3 model and propose a multi-scale cross-attention difference siamese network (MC-DiSNet). The powerful visual representation capability of the DINOv3 backbone enables robust and rich feature extraction from bi-temporal remote sensing images. The multi-scale cross-attention mechanism allows for precise localization of subtle semantic changes, while the difference siamese structure enhances inter-class feature discrimination, enabling fine-grained semantic change detection. Furthermore, a simple yet powerful lightweight decoder is designed to generate clear detection maps while maintaining high efficiency. We also release a new Gaza-change dataset containing high-resolution satellite image pairs from 2023-2024 with pixel-level semantic change annotations. It is worth emphasizing that our annotations only include semantic pixels of changed areas. Unlike conventional semantic change detection (SCD), our approach eliminates the need for large-scale semantic annotations of bi-temporal images, instead focusing directly on the changed regions, which terms multi-class change detection (MCD). We evaluated our method on the Gaza-Change and two classical datasets: SECOND and Landsat-SCD datasets. Experimental results demonstrate that our proposed approach effectively addresses the MCD task, and its outstanding performance paves the way for practical applications in rapid damage assessment across conflict zones.
@@ 1 Introduction
To address the aforementioned challenges, we introduce the multi-class change detection (MSD). Distinct from conventional semantic change detection (SCD), MCD eliminates the need for annotating entire semantic regions, instead focusing solely on change masks. This framework represents a direct extension of binary change detection (BCD). While significantly reducing annotation difficulty and time requirements, this new paradigm consequently increases the challenge of limited target region proportions. This necessitates substantial improvements in the model’s capability to extract features from small target areas.
@@
Figure 1: Evolution of change detection paradigms: (a) Binary Change Detection (BCD), (b) Semantic Change Detection (SCD), and (c) Multi-Class Change Detection (MCD).
@@
Based on the above analysis, we identify four major challenges in framing conflict-induced damage assessment as a MCD task: (1) Inherent data scarcity: Limited by the geographical extent of conflict zones and the number of destroyed areas available for training. (2) Small target regions: MCD focuses exclusively on damaged areas, resulting in minimal semantic region coverage. (3) Subtle and ambiguous changes: Infrastructure damage in conflict zones varies significantly in severity and extent, particularly making minor damage difficult to detect. (4) High inter-class similarity: Different facility categories within the same region may share similar characteristics, making fine-grained damage assessment particularly challenging for semantic change detection.
@@
To bridge this gap, we draw inspiration from the recent success of foundational vision models. We argue that leveraging their rich, pre-trained representations is key to overcoming data scarcity and recognizing subtle semantic changes. In this paper, we propose a novel DINOv3-driven siamese network for MCD. Specifically, we adopt the DINOv3 simeoni2025dinov3 model pre-trained on satellite data, with ConvNeXt liu2022convnet as its main backbone architecture, which helps reduce the distribution discrepancy between the pre-training data and the actual application data. Then, we propose a multi-scale attention mechanism to extract and enhance features at different levels, aiming to capture the subtle and ambiguous change features of infrastructure damage. Furthermore, we perform an absolute value differential operation on the obtained semantic-rich feature maps to increase inter-class feature differences. Finally, a carefully designed decoder network with attention enhancement is used to generate clear semantic change detection maps. We also release a building semantic change detection dataset of the Gaza area from 2023 to 2024. As shown in Figure 2, we present panoramic remote sensing images of the Gaza Strip captured by satellites. To the best of our knowledge, this is the first remote sensing semantic change detection study focused on conflict area assessment, laying a foundation for future research in related fields. In summary, our work makes the following key contributions:
@@
  • We introduce the multi-class change detection (MCD) paradigm for damage assessment that fundamentally shifts from exhaustive bi-temporal semantic annotation to focused labeling of changed semantic regions. This strategic simplification significantly reduces annotation complexity and human labor.
  • Extensive experiments show that our method achieves state-of-the-art performance not only on our proposed Gaza dataset but also on the public benchmarks, SECOND yang2021asymmetric and Landsat-SCD yuan2022transformer, demonstrating its superior robustness and generalization capability.
@@ 3.2 Enhanced DINOv3 Backbone
To maximize the feature extraction capability of the DINOv3 backbone while balancing computational cost and accuracy for the multi-class change detection (MCD) task, we introduce an enhanced DINOv3 encoder. This encoder keeps the original ConvNeXt-Tiny weights frozen and enriches multi-scale features C3-C5 through three lightweight plug-ins. First, a bottleneck adapter is inserted after each MLP block, comprising a down-projection, GELU activation lee2023mathematical, and an up-projection branch with a reduction ratio. Its forward pass is formulated as:
@@ 3.3 Multi-Scale Cross-Attention Difference Module
Next, the concatenated tensor $\mathbf{F}_{\mathrm{cat}}$ == $\mathbf{C}_{\mathrm{fused}}^{t_{1}};\mathbf{C}_{\mathrm{fused}}^{t_{2}}]$ is fed to a lightweight stack of $3\times 3$ conv, BN, GELU and $1\times 1$ conv to model non-linear temporal discrepancies and produce an adaptive difference feature $\mathbf{D}_{\mathrm{ada}}$ (diff_ada in Figure 3).
Finally, the direct difference $\mathbf{D}_{\mathrm{dir}}=|\mathbf{C}^{t_{1}}-\mathbf{C}^{t_{2}}|$ (diff_abs in Figure 3) and the adaptive difference $\mathbf{D}_{\mathrm{ada}}$ are concatenated and compressed by $1\times 1$ conv, BN and GELU, yielding the multi-scale change feature diff_aggergator:
@@
3.4 Decoder
To refine the coarse change cues produced by the MSCAD module, we design a enhancement decoder, which contains two lightweight yet effective sub-blocks. First, a three-layer depthwise-separable residual stack (Figure 3) enriches spatial context via
@@ 3.4 Decoder
$\mathbf{F}_{\mathrm{out}}=\mathbf{G}\odot\mathbf{Y},$ (8)
@@ 4.1.1 Datasets Description
Landsat-SCD yuan2022transformer provides ten change types and is constructed based on Landsat-like images acquired between 1990 and 2020 in Tumushuke, Xinjiang. The region is located along the Belt and Road Economic Belt, adjacent to the Taklimakan Desert, and features a fragile ecological environment. The data source consists of Landsat series imagery with a spatial resolution of 30 meters, which offers relatively high spatiotemporal and spectral resolution as well as good data accessibility.
Table 1 summarizes the specific information of the three aforementioned datasets. It is important to emphasize that while both the Gaza-Change, SECOND, and Landsat-SCD datasets are annotated with semantic labels. To validate the MCD task, the labels in the SECOND and Landsat-SCD dataset were also processed to retain semantic information exclusively for the changed areas.
Table 1: Main information of three datasets.
@@ 4.1.3 Experimental Details
All compared methods are described as follows:
@@
4.2 Main Results
Result of Caza-change dataset. We report the results achieved by the model checkpoint that performs best on the test set. Table 2 summarizes the quantitative comparison between MC-DiSNet and nine representative methods.
@@ 4.2 Main Results
Method Trainable Params (M) P(%) R(%) mIoU(%) F1(%) Best Epoch
USSFCNET 1.5 46.66 48.75 34.43 47.20 51
BIT 11.9 57.45 54.90 42.13 56.15 65
SNUNet 12.0 51.44 51.42 36.50 50.35 49
DDLNET 13.8 56.61 59.96 42.23 57.22 34
STNet 14.6 59.15 54.96 42.81 56.64 53
CDxLstm 16.2 56.96 47.29 36.84 51.25 83
ChangeFormer 41.1 59.06 46.52 36.62 50.50 88
Rsmamba 52.0 58.43 42.81 36.45 49.40 66
LGPNet 71.0 55.69 55.69 48.64 54.64 64
Ours 18.2 86.10 60.98 55.16 69.25 29
Results of SECOND and Landsat-SCD dataset. To further verify the effectiveness and generalizability of the proposed method, we conduct experiments on the two extra SCD datasets: SECOND and Landsat-SCD. Quantitative results are summarised in Table 3.
Obviously, our method achieves an mIoU of 27.74% and an F1-score of 40.59%, surpassing the strongest competitor (BIT) by 6.87 percentage points in mIoU and 8.39 percentage points in F1-score. On the large-scale Landsat-SCD dataset, our method achieves a mIoU of 68.79% and an F1-score of 80.69%, significantly outperforming the previous state-of-the-art LGPNet (61.48% mIoU and 75.45% F1). This represents notable absolute gains of +7.31% in mIoU and +5.24% in F1, corresponding to relative improvements of 11.9% and 6.9%, respectively. This consistent superiority on two external datasets demonstrates that the proposed modules do not overfit the primary training data and generalize effectively to other remote sensing scenarios with varying imaging conditions and annotation granularity.
Table 3: Performance comparison on SECOND and Landsat-SCD datasets.
Method SECOND Landsat-SCD
mIoU (%) F1 (%) mIoU (%) F1 (%)
BIT 20.87 32.20 60.13 74.33
SNUNet 12.88 20.85 42.07 56.28
USSFCNet 12.03 17.80 41.05 55.14
ChangeFormer 15.48 28.30 57.09 71.45
CDxLstm 15.19 27.25 47.93 62.89
RSMamba 12.54 20.09 55.41 70.00
LGPNet 12.42 19.37 61.48 75.45
Ours 27.74 40.59 68.79 80.69
Figure 7: Example results on Landsat-SCD dataset.
Figure 8: Qualitative comparison on the Landsat-SCD dataset, with color coding: true positives (white), true negatives (black), false positives (red), and false negatives (green).
We also present visual detection results on the Landsat-SCD dataset in Figure 7 and Figure 8. As shown in the third row of Figure 7, MC-DiSNet produces a continuous and uniformly wide delineation of the narrow, curved change belt, whereas results from other methods exhibit fragmentation or over-expansion. The first row of Figure 8 demonstrates that our result aligns closely with the ground truth (GT) without spurious speckles, while other methods show extensive green false negatives (FN). However, it is important to note that while our method significantly reduces under-detection, it introduces a certain amount of false positives (red pixels), leading to an overall reddish hue in some areas. This observation underscores a broader challenge: effective change detection models often require targeted adaptation to the specific characteristics of different remote sensing datasets.
Overall, As demonstrated on our own dataset as well as on the additional benchmarks, the proposed approach consistently delivers state-of-the-art results not only on small-scale sets (hundreds of image pairs) but also on large-scale collections (tens of thousands of samples). This robust performance across dramatically different data volumes indicates strong generalization capability and practical deployment potential for diverse remote-sensing scenarios.
@@
Module MS-att diff-ada diff-agg att P(%) R(%) mIoU (%) F1 (%)
Encoder 89.37 54.35 50.28 64.99
79.26 49.20 44.25 58.36
78.77 50.50 46.10 59.09
Decoder 83.39 60.24 53.81 67.26
Base 86.10 60.98 55.16 69.25
Table 5: Comparison with different backbones in encoder of our proposed MC-DiSNet on the Gaza-change dataset.
Method P(%) R(%) mIoU(%) F1(%) Params (M) Trainable Params (M)
Dinov3-tiny (ours) 86.10 60.98 55.16 69.25 29.09 1.78
Dinov3-smal1 77.52 65.35 62.03 70.71 53.06 3.25
ResNet-18 51.56 40.13 32.45 42.99 14.95 0.68
ResNet-50 59.37 39.07 34.03 44.48 25.76 1.55
SAM-base 90.61 59.95 56.46 71.65 95.39 1.36
Swinformer-tiny 73.71 63.47 58.05 67.92 29.58 1.62
@@
Table 7: Quantitative results of our MC-DiSNet on the proposed Gaza-change dataset under the MCD task, reporting per-category performance metrics for all six change types.
@@
4.3 Ablation Study
4.3.1 Module-level Ablation
We conduct ablation studies on key components of MC-DiSNet: (a) the multi-scale attention (MS-att) in the encoder, (b) the two feature fusion modules, diff_Ada and diff_agg, and (c) the attention mechanism (Att) in the decoder. Table 4 reports the comparison results.
To systematically validate the necessity and synergistic effects of the core components in MC-DiSNet, we conducted module ablation experiments on the validation set, with results presented in Table 4. The full model with all components (Base) achieved the best performance (55.16% mIoU, 69.25% F1). Removing components individually led to performance degradation to varying degrees: Removing the multi-scale attention (MS-att) caused mIoU to drop by 4.88% and F1 by 4.26%, indicating its crucial role in multi-scale discriminative feature selection. Removing the decoder attention (att) reduced mIoU to 53.81% and lowered the recall rate, confirming its core function in refining the final change map. Removing diff_agg resulted in one of the most severe performance drops (mIoU 46.10%, F1 59.09%), highlighting the irreplaceability of cross-level fusion of difference features. Finally, completely removing diff_ada led to the largest degradation (mIoU 44.25%, F1 58.36%), demonstrating its key role in suppressing false-change noise.
[…]
4.3.2 Parameter-level Ablation
Table 6 compares different combinations of rank $r$ and scaling factor $\alpha$. As shown in Table.6, the best trade-off between accuracy and parameter budget is achieved at r=24 and $\alpha$=48, where the model reaches 69.25% F1 and 55.16% mIoU with only 0.97 M additional trainable weights.
5 Discussion
Second, we introduce the MCD task that directly annotates changed areas rather than separately labeling bi-temporal semantic regions. This approach effectively extends traditional Binary Change Detection (BCD) to semantic change analysis while avoiding error accumulation from intermediate processing stages. The MCD framework substantially reduces both annotation workload and complexity, enabling rapid damage assessment in conflict scenarios.
@@ 5 Discussion
These findings highlight both the promise and challenges of automated damage assessment in conflict zones, while demonstrating the potential of the MCD paradigm and MC-DiSNet architecture to advance the field of humanitarian remote sensing.